There are 3 models that generate 3 below predicted values for CAD
Model1: $1.00
Model2: $1.112
Model3: $0.982
You assign the following weight for these 3 models
Model1: 50%
Model2: 20%
Model3: 30%
What is the mix forecast value for GBP?
Answer #1

Calculation of mix forcast value:

Mix forcast value= (1*0.50)+(1.112*0.20)+(0.982*0.30)

Mix forcast value= 0.50+0.2224+0.2946= $1.017

Mix forcast value is $1.017